# A Nuclear Interferometer for Ultra-Light Dark Matter Detection

### v1

@@ III.2 Neutral Atoms
One could alternatively conceive a nuclear interferometer based upon the ground - isomeric excited state transition in neutral 229Th atoms. Whilst this allows for significantly greater effective flux rates and thus much lower quantum projection noise, there are additional complexities relating to the short lifetime of the excited state which need to carefully considered. Compared to clock interferometry, in the case of a broad transition the experimental sensitivity is negatively impacted by spontaneous decay occurring both during periods of free propagation in the excited state and during atom-laser interactions. Each of these effects can be accounted for by the introduction of an appropriate correction factor as we shall now detail.
@@
Eq. 27 thus implies that there is a minimum $\pi$-pulse duration for each $l$, which, for the transition in question is sizeable relative to the excited state lifetime. It thus becomes important to account for the impact of spontaneous decay over the duration of the $\pi$-pulse. This can be formally achieved by explicitly including spontaneous decay in the optical Bloch equations which describe the behaviour of a 2-level system interacting with an electromagnetic field [96, 97]. The result, which we detail explicitly in Appendix B, is a reduction in the probability that an applied $\pi$-pulse results in population inversion from its ideal value of unity according to the value of the ratio $\Omega/\Gamma_{a}$.
@@
where $P_{k}$ denotes the maximum efficiency of a $\pi$-pulse that addresses a single interferometer arm at an $l=k$ velocity separation, obtained by evaluating Eq. 39 at the maximum ratio of $\Omega/\Gamma_{a}$ according to Eqs. 27 and 29.
@@
In our sensitivity forecasts for this scenario we consider three distinct operating configurations, characterised by the parameters $T,$ $L$ and $n$, the interferometer separation $\Delta z$ and the intrinsic phase noise $\sqrt{\tilde{S}_{\textnormal{n}}}$ that arises from the number of atoms launched per shot, representative of an ‘initial’, ‘intermediate’ and more ‘advanced’ experimental design. Whilst we specify our configurations using $\sqrt{\tilde{S}_{\textnormal{n}}}$, we emphasise that it is $S_{\textnormal{n}}$, as defined in Eq. 31, which actually enters the sensitivity calculations in order to account for the additional loss of coherence/atom number due to spontaneous decay. Some justification of these values, which we detail in Tab. 2, is in order. As explained above, due to the substantial impact of spontaneous decay on the $\pi$-pulse efficiency we are restricted to an $n=2$ LMT order. According to Eq. 23, the greatest sensitivity will then be reached at a baseline of $\sim$ 750 m which we thus deploy for our advanced configuration. In order to take advantage of existing and planned experimental infrastructure, for the initial and intermediate configurations we specifically choose values for $L$ that match the baselines that are, or will be, deployed in the prototype iterations of proposed very-long-baseline clock interferometers based on the 87Sr clock transition.
@@
As per Eq. 19, in the shot-noise limit the phase sensitivity is essentially controlled at the design level by the time-averaged atomic flux ($N_{\textnormal{s}}/\Delta t$). Given that atom interferometry relies upon the ensemble of atoms being sufficiently cold such that the fraction of atoms that are addressed with each laser pulse is not significantly degraded by the thermal expansion of the cloud over the course of the interferometer sequence, it is laser cooling techniques which ultimately limit the achievable flux rates. As discussed in detail in Refs. [46, 98], the percentage of atoms which interact with each laser pulse follows from a convolution of the spatial and velocity distributions of the ensemble of atoms with the laser pulse Fourier transform and spatial profile of the beam. In our forecasts we assume that the laser beam waist in use is sufficient to address every atom in every interaction. Since the clouds of atoms expand over time as a result of their thermal velocity, the initial temperature to which the atoms are cooled thus needs to be sufficiently low such that the cloud size remains appropriate for the beam waist in use. A greater understanding of the cooling temperatures and laser properties that may be required to achieve this could be developed through a more thorough analysis of the convolution integral. Assuming the usage of a similar beam waist, the requirements on the 3D momentum distribution of an ensemble of thorium atoms are likely to be of similar order of magnitude to those of more typical atom interferometry sources such as strontium. We thus inform our forecasts of potential fluxes with the initial and more ambitions projections discussed in the context of conventional atom interferometers (see e.g. Ref. [49]).
@@
A further issue which one needs to remain mindful of is that in neutral 229Th, the nuclear excitation energy lies above the first ionisation potential. [^12] Note that this is not the case in ionic form. Whilst a first-principles evaluation of the photo-ionisation and photo-excitation rates of this system lies beyond the scope of this work, we provide an estimation as to the potential implications of this in Appendix C. We find that whilst photo-ionisation is likely to make working with unpolarised atoms infeasible, if the clouds are polarised it may be possible via a careful choice of linear laser polarisation to greatly suppress the photo-ionisation rate by forbidding the dominant photo-ionisation channels. The projections shown in Sec. IV assume the implementation of such a practice such that atom loss due to ionisation is negligible.
@@ IV Discussion
In Figs. 2, 3 and 4, we plot the couplings $d_{e}$, $d_{g}$ and $d_{\hat{m}}$ that could be detected at a signal-to-noise ratio (SNR) of 1 with the single-ion interferometer designs specified in Tab. 1 following a measurement period of $T_{\textnormal{int}}=10^{8}$ s, assuming that the coupling in question dominates $d_{\phi}$. In Figs. 5, 6 and 7, we present the reach of the neutral atomic interferometer configurations of Tab. 2, once again taking $T_{\textnormal{int}}=10^{8}$ s and assuming that the clouds of atoms have been appropriately polarised such that photo-ionisation is negligible (see Appendix C). For clarity, our plots show the power-averaged envelope of the experimental response using the approximation $\sin x$ = min($x$, $1/\sqrt{2}$) as in Ref. [39].
Figure 2: Shot-noise limited projected sensitivity of the single-ion nuclear interferometer configurations detailed in Tab. 1 to ULDM with a linear scalar coupling to photons. The curves show the minimum coupling $d_{e}$ that could be detected at a SNR = 1 following an experimental campaign of duration $T_{\textnormal{int}}=10^{8}$s. Also shown are existing bounds and forecast sensitivities of other experiments as detailed in the main text. The parameter space motivated from a naturalness perspective as described in the text is shown in red.
@@
In the case of the coupling to photons $d_{e}$, we also show limits from the AURIGA detector [107], the frequency comparison of the electric octupole and the electric quadrupole clock transitions in single-ion ytterbium (Yb+ E3/E2), the comparison of the Yb+ E3 transition to the strontium $S\to P$ clock transition (Yb+ E3/Sr)  [37], the comparison of a strontium optical clock with a silicon cavity (Sr/Si) [108] and the combined constraints of several optical clock comparisons (Al+/Hg+, Al+/Yb, Yb/Sr) from the Boulder atomic clock optical network [36]. Forecasts for the sensitivities of proposed compact mechanical resonators based on superfluid helium, sapphire cylinders and sapphire pillars [34] are also shown, as are the prospective sensitivities of very-long-baseline atom interferometers operating on the 87Sr atomic clock transition. We show forecasts for terrestrial configurations (TVLBAIs) with 10 m, 100 m and 1 km scale baselines in addition to a potential space-based realisation (SB-VLBAI), assuming the experimental parameters listed in Tab. 3. Further details on the calculation of these estimates, including a justification of the parameters used, can be found in Appendix D.
@@
Figure 5: Shot-noise limited projected sensitivity of the nuclear interferometer configurations deploying neutral atoms detailed in Tab. 2 to ULDM with a linear scalar coupling to photons. The curves show the minimum coupling $d_{e}$ that could be detected at a SNR = 1 following an experimental campaign of duration $T_{\textnormal{int}}=10^{8}$ s and assuming that photo-ionisation has been successfully suppressed. Also shown are existing bounds and forecast sensitivities of other experiments as detailed in the main text. The parameter space motivated from a naturalness perspective as described in the text is shown in red.
@@
As a further point of note, we emphasise that our forecasts assume the experiment to be shot-noise limited at all frequencies. In practice, for terrestrial experiments, gravitational gradient noise (GGN) arising from mass fluctuations around the experiment [114, 115] is thought to dominate at sub-Hz frequencies [116, 117, 118]. The modelling and mitigation of GGN at terrestrial atom interferometers constitutes an active and involved area of research (see e.g. Ref. [119]) which is beyond the scope of this work. As such, we do not attempt to take the impact of GGN into account here but caution the reader that below $\sim$1 Hz, our estimations for the terrestrial nuclear interferometer configurations (and indeed those of the TVLBAIs) should be read as extrapolations of the reach that could be attained if GGN could be sufficiently mitigated such that the experiment remains shot-noise limited. As demonstrated in Ref. [119], the use of multiple interferometers positioned along the baseline could be promising in this respect, with the potential to reduce the impact of GGN in ULDM searches by up to two orders of magnitude.

### v2

@@ III.2 Neutral Atoms
One could alternatively conceive a nuclear interferometer based upon the ground - isomeric excited state transition in neutral 229Th atoms. Whilst this allows for significantly greater effective flux rates and thus much lower quantum projection noise, there are additional complexities relating to the short lifetime of the excited state which need to be carefully considered. Compared to clock interferometry, in the case of a broad transition the experimental sensitivity is negatively impacted by spontaneous decay occurring both during periods of free propagation in the excited state and during atom-laser interactions. Each of these effects can be accounted for by the introduction of an appropriate correction factor as we shall now detail.
@@
Eq. 29 therefore implies that there is a minimum $\pi$-pulse duration for each $l$, which, for the transition in question is sizeable relative to the excited state lifetime. Hence it becomes important to account for the impact of spontaneous decay over the duration of the $\pi$-pulse. This can be formally achieved by explicitly including spontaneous decay in the optical Bloch equations which describe the behaviour of a 2-level system interacting with an electromagnetic field [117, 118]. The result, which we detail explicitly in Appendix B, is a reduction in the probability that an applied $\pi$-pulse results in population inversion from its ideal value of unity according to the value of the ratio $\Omega/\Gamma_{a}$.
@@
where $P_{k}$ denotes the maximum efficiency of a $\pi$-pulse that addresses a single interferometer arm at an $l=k$ velocity separation, obtained by evaluating Eq. 42 at the maximum ratio of $\Omega/\Gamma_{a}$ according to Eqs. 29 and 31.
@@
In our sensitivity forecasts for this scenario, we consider three distinct operating configurations, characterised by the parameters $T,$ $L$ and $n$, the interferometer separation $\Delta z$ and the intrinsic phase noise $\sqrt{\tilde{S}_{\textnormal{n}}}$ that arises from the number of atoms launched per shot, representative of an ‘initial’, ‘intermediate’ and more ‘advanced’ experimental design. Whilst we specify our configurations using $\sqrt{\tilde{S}_{\textnormal{n}}}$, we emphasise that it is $S_{\textnormal{n}}$, as defined in Eq. 33, which actually enters the sensitivity calculations in order to account for the additional loss of coherence/atom number due to spontaneous decay. Some justification of these values, which we detail in Tab. 2, is in order. As explained above, due to the substantial impact of spontaneous decay on the $\pi$-pulse efficiency we are restricted to an $n=2$ LMT order. According to Eq. 25, the greatest sensitivity will then be reached at a baseline of $\sim$ 750 m which we thus deploy for our advanced configuration. In order to take advantage of existing and planned experimental infrastructure, for the initial and intermediate configurations we specifically choose values for $L$ that match the baselines that are, or will be, deployed in the prototype iterations of proposed very-long-baseline clock interferometers based on the 87Sr clock transition.
@@
As per Eq. 19, in the shot-noise limit the phase sensitivity is essentially controlled at the design level by the time-averaged atomic flux ($N_{\textnormal{s}}/\Delta t$). Given that atom interferometry relies upon the ensemble of atoms being sufficiently cold such that the fraction of atoms that are addressed with each laser pulse is not significantly degraded by the thermal expansion of the cloud over the course of the interferometer sequence, it is laser cooling techniques which ultimately limit the achievable flux rates. As discussed in detail in Refs. [49, 119], the percentage of atoms which interact with each laser pulse follows from a convolution of the spatial and velocity distributions of the ensemble of atoms with the laser pulse Fourier transform and spatial profile of the beam. In our forecasts we assume that the laser beam waist in use is sufficient to address every atom in every interaction. Since the clouds of atoms expand over time as a result of their thermal velocity, the initial temperature to which the atoms are cooled thus needs to be sufficiently low such that the cloud size remains appropriate for the beam waist in use. A greater understanding of the cooling temperatures and laser properties that may be required to achieve this could be developed through a more thorough analysis of the convolution integral. Assuming the usage of a similar beam waist, the requirements on the 3D momentum distribution of an ensemble of thorium atoms are likely to be of similar order of magnitude to those of more typical atom interferometry sources such as strontium (i.e. at the nK - pK level [99, 49]). We thus inform our forecasts of potential fluxes with the initial and more ambitions projections discussed in the context of conventional atom interferometers (see e.g. Ref. [52]).
@@
A second significant challenge of working with neutral 229Th is that the 8.3 eV nuclear excitation energy is greater than the first ionisation potential of the atom [^16] Note that this is not the case for ionic forms. which is 6.3 eV. With the interferometry schemes proposed in this work being reliant on the successful completion of a precise sequence of coherent, laser-induced nuclear excitations and de-excitations, it is necessary to explore the extent to which photo-ionisation would compete with these processes, depleting the number of atoms at the point of detection and reducing the sensitivity. Whilst a first-principles evaluation of the photo-ionisation and photo-excitation rates of this system lies beyond the scope of this work, we offer a rough preliminary assessment here by comparing estimations of the photo-ionisation and nuclear photo-excitation cross-sections of atomic 229Th in literature.
From the ground electronic configuration, a laser resonant with the isomeric transition is sufficiently energetic to ionise the atom via removal of electrons from either the 6d or 7s electronic shells. At leading order, these processes proceed via allowed (E1) electric dipole transitions. The cross-sections for E1 photo-ionisation of the 6d and 7s electrons are computed numerically in Ref. [122] using the flexible atomic code (FAC) based upon the relativistic configuration interaction method [123]. At an energy of $\sim$ 8 eV these cross-sections are of order 10-17 cm2 [122]. In the low saturation limit of moderate laser intensities and a laser bandwidth that exceeds the natural linewidth of the transition, the resonance nuclear photo-excitation cross-section is given by [124]
where $\Gamma_{\gamma}\sim 10^{-9}\hskip 5.0pt\Gamma_{a}$ is the radiative decay rate and $\lambda_{0}$ is the transition wavelength. Setting $\Gamma_{l}$ to $\Gamma_{a}$, its minimum value consistent with the regime for which this equation is valid, gives $\sigma_{\textnormal{ex.}}\sim 10^{-20}$ cm2, approximately three orders of magnitude below the photo-ionisation rate. Whilst, as stated in Appendix B, our sensitivity forecasts assume the laser bandwidth to be narrow relative to the natural linewidth to reduce atom loss from spontaneous decay, here we are just interested in an order of magnitude estimate of the nuclear excitation rate in order to assess the extent to which photo-ionisation is likely to be a problem. Given that the nuclear excitation rate should be greatest in the low saturation limit, using this result provides a conservative assessment of the potential impact of photo-ionisation on the interferometry sequence and should thus suffice here.
These preliminary estimates indicate that, unless some means with which to suppress ionisation can be developed, performing interferometry with neutral thorium is unlikely to be feasible due to a significant depletion of the number of atoms remaining coherently in the cloud at detection. Given that the cross-sections of higher order (e.g. first forbidden magnetic dipole M1) ionisation processes are, according to the computations in Ref. [122], approximately 5 orders of magnitude smaller than estimations of the nuclear photo-excitation rate, it should suffice to suppress just the two leading order E1 ionisation channels identified previously. To this end it may be possible to exploit arguments based on transition selection rules to achieve some degree of suppression by working with polarised clouds of atoms and making use of a careful choice of laser polarisation. A number of experimental studies have observed that the total photo-ionisation cross-section of polarised atoms depends on the relative orientation of the atom and laser polarisation (see Refs. [125, 126, 127, 128]). However, determining the optimal configuration, in addition to assessing the extent of suppression that may result, demands a full atomic calculation, and falls beyond the remits of this study. It may be that such a practice is insufficient to reach the desired 3 order of magnitude suppression on its own, and that it is necessary to look to other more active methods to achieve this. Various such techniques have been the subject of both theoretical and experimental attention in literature (see Refs. [129, 130, 131, 132, 133, 134]) and have been shown capable of suppressing photo-ionisation by several orders of magnitude [135]. The assessment of whether or not such methods (or indeed others) could be adapted to suppress photo-ionisation in 229Th in a way that is compatible with the proposed interferometry sequence lies beyond the scope of this work.
In light of this, in the forecasts presented in Sec. IV of this work we show the reach that could be achieved should it be possible to develop appropriate techniques to suppress photo-ionisation to negligible levels. In this way our estimates serve as motivation from a fundamental physics perspective for a dedicated programme aimed at establishing the feasibility of, and subsequently developing, appropriate atom stabilising technology for 229Th.
Whereas the leading source of noise other than shot-noise affecting the aforementioned single-ion proposal was magnetic field noise, given both the electrical neutrality of atoms and the insensitivity of the nuclear clock transition to perturbations from external fields, such fluctuations are much less of a concern here. The attainment of shot-noise limited sensitivity in this case is instead challenged by so-called gravitational gradient noise (GGN) [^17] Whilst GGN similarly affects the ionic proposal previously discussed, it is substantially smaller than the high shot-noise resulting from single ion shots and thus does not limit the sensitivity. which arises due to mass density fluctuations of the ground and atmosphere around the experiment [136, 137]. These fluctuations perturb the local gravitational potential experienced by the clouds of atoms, generating spurious gradiometer phase shifts. This type of noise cannot be shielded and has been identified as a major challenge for conventional terrestrial very-long-baseline atom interferometry proposals [52]. For typical 100 m- and km-scale instruments, GGN is to expected surpass the shot-noise at $\sim$ sub-Hz frequencies, leading to a significant loss of sensitivity [138, 139, 140, 101, 141].
The modelling and mitigation of GGN at terrestrial atom interferometers constitutes an active and involved area of research (see e.g. Ref. [141]) which depends both on details of the interferometry sequence and the local seismic environment of the experiment. Given this latter consideration, we do not attempt to perform any first-principles calculations here, but instead estimate the potential impact of GGN on the atom-based nuclear interferometer designs specified in Tab. 2 by adapting existing analyses for standard long-baseline atom interferometers presented in the literature to account for differences in the interferometry sequence and instrument design.
The GGN spectral density for both a 100 m- and a km-scale atom interferometer with the design specifications detailed in Tab. 3 were calculated in Ref. [141] in the $1010^{-3}-$ 10 Hz frequency band by modelling the seismic field as an incoherent superposition of monochromatic plane waves propagating isotropically at the Earth’s surface. Results for both a representative low and high noise seismic environment were presented, based respectively on the empirically driven Peterson new low and high noise models [142].
At low frequencies, the GGN spectral noise density, $S_{\textnormal{ggn}}$, scales approximately with the interferometer design parameters as $\sim n^{2}L^{2}T^{4}$ [141]. To obtain a conservative estimate the impact of GGN on the nuclear interferometer configurations considered in this work, we re-scale the high noise $S_{\textnormal{ggn}}$ presented in Ref. [141] according to the parameters of our instruments and locate the frequency at which this becomes dominant over the shot-noise. [^18] We emphasise that it is $S_{n}$, the shot-noise spectral density accounting for spontaneous decay which should be used for this comparison. Due to both the substantially lower $n$ and the increased shot-noise (due to the corrections for spontaneous decay) of the intermediate and advanced nuclear interferometer designs proposed here, we find that they remain shot-noise limited down to lower frequencies: $\sim 10^{-3}$ Hz and $\sim 10^{-2}$ Hz respectively, compared to the $\sim$ Hz threshold of the 100 m and km instruments of Tab 3.
The analysis of Ref. [141] found that the sensitivity of the 10 m instrument defined in Tab. 3 was not impacted by GGN in the frequency band considered in either the low or high noise scenarios. It is therefore safe to assume that our 10 m initial nuclear interferometer would also remain in the shot-noise limited regime above $10^{-3}$ Hz, however a full calculation of the GGN spectral density at lower frequencies would be required to obtain a more precise handle on the frequency down to which this is the case. For this reason, when we delineate the frequencies at which different instruments may become GGN-limited on our plots in Sec. IV, we conservatively take a $10^{-3}$ Hz threshold for the 10 m instruments of both Tab. 2 and Tab. 3.
Given the potential impact of GGN at terrestrial atom interferometers, various passive mitigation techniques, which may also be applicable to nuclear interferometry, are under active investigation. In particular it was shown in Ref.  [141] that multigradiometer configurations in which multiple interferometers are positioned along the same baseline, could greatly reduce the impact of GGN on ULDM searches and recover significant regions of parameter space accessible to shot-noise limited instruments. A carefully selected site location with a quiet seismic environment (i.e. approaching the Peterson low noise as opposed to high noise model) may also offer the opportunity to regain up to three orders of magnitude in sensitivity [52].
@@ IV Discussion
In Figs. 2, 3 and 4, we plot the couplings $d_{e}$, $d_{g}$ and $d_{\hat{m}}$ that could be detected at a signal-to-noise ratio (SNR) of 1 with the single-ion interferometer designs specified in Tab. 1 following a measurement period of $T_{\textnormal{int}}=10^{8}$ s, assuming that the coupling in question dominates $d_{\phi}$. In Figs. 5, 6 and 7, we present the shot-noise limited reach of the neutral atomic interferometer configurations of Tab. 2, once again taking $T_{\textnormal{int}}=10^{8}$ s and assuming the suppression of photo-ionisation to negligible levels. For clarity, our plots show the power-averaged envelope of the experimental response using the approximation $\sin x$ = min($x$, $1/\sqrt{2}$) as in Ref. [40].
We emphasise that our forecasts assume the nuclear interferometer sensitivity to be shot-noise limited over the entire frequency band of the experiment. As discussed in Sec. III, there may however be other sizeable noise sources which could become dominant. In order to reflect this, we indicate on our plots the frequencies, as determined by the analysis in Sec. III, at which the experiment may depart from the shot-noise limited regime due to either magnetic field noise (in the case of ions) or GGN (in the case of atoms) with the use of a dotted, as opposed to a solid, line. We highlight that these thresholds assume that no attempts to mitigate either noise source either actively or passively are made. Advancements in large-scale magnetic shielding technology, a well-chosen site location and close environmental monitoring could all help to extend the shot-noise limited regime into lower frequencies and to recover regions of the impacted parameter space.
Figure 2: Shot-noise limited projected sensitivity of the single-ion nuclear interferometer configurations detailed in Tab. 1 to ULDM with a linear scalar coupling to photons. The curves show the minimum coupling $d_{e}$ that could be detected at a SNR = 1 following an experimental campaign of duration $T_{\textnormal{int}}=10^{8}$s. The dotted portion of the curve indicates where the sensitivity may be adversely impacted by magnetic field noise according to the discussion in Sec. III. Also shown are existing bounds and forecast sensitivities of other experiments as detailed in the main text. The parameter space motivated from a naturalness perspective as described in the text is shown in red.
@@
In the case of the coupling to photons $d_{e}$, we also show limits from the AURIGA detector [149], the frequency comparison of the electric octupole and the electric quadrupole clock transitions in single-ion ytterbium (Yb+ E3/E2), the comparison of the Yb+ E3 transition to the strontium $S\to P$ clock transition (Yb+ E3/Sr)  [37], the comparison of a strontium optical clock with a silicon cavity (Sr/Si) [150] and the combined constraints of several optical clock comparisons (Al+/Hg+, Al+/Yb, Yb/Sr) from the Boulder atomic clock optical network [36]. Forecasts for the sensitivities of proposed compact mechanical resonators based on superfluid helium, sapphire cylinders and sapphire pillars [34] are also shown, as are the prospective shot-noise limited sensitivities of very-long-baseline atom interferometers operating on the 87Sr atomic clock transition. We show forecasts for terrestrial configurations (TVLBAIs) with 10 m, 100 m and 1 km scale baselines in addition to a potential space-based realisation (SB-VLBAI), assuming the experimental parameters listed in Tab. 3. Further details on the calculation of these estimates, including a justification of the parameters used, can be found in Appendix C. As for the atom-based nuclear interferometers we indicate the frequencies at which the sensitivities of these experiments are likely to be impacted by GGN, assuming both a high seismic environment and that no mitigation strategies are applied, according to the analysis in Ref. [141], by switching from a dashed to a dotted line.
@@
Figure 5: Shot-noise limited projected sensitivity of the nuclear interferometer configurations deploying neutral atoms detailed in Tab. 2 to ULDM with a linear scalar coupling to photons. The curves show the minimum coupling $d_{e}$ that could be detected at a SNR = 1 following an experimental campaign of duration $T_{\textnormal{int}}=10^{8}$ s and assuming that photo-ionisation has been successfully suppressed. The dotted portion of the curve indicates where the sensitivity may be adversely impacted by GGN according to the discussion in Sec. III. Also shown are existing bounds and forecast sensitivities of other experiments as detailed in the main text. The parameter space motivated from a naturalness perspective as described in the text is shown in red.
@@
A further interesting line of enquiry could be to investigate the possibility of binding 229Th into a molecule. Not only would it be necessary to verify that the nuclear transition, and its enhanced sensitivity to the variation of fundamental constants remains intact in such a system, but that both the dissociation and ionisation energies of the molecules lie above the nuclear transition energy. To fully harness the benefits that working with a neutral species presents in terms of the shot-noise, it would be desirable for the excited nuclear state to be long lived, avoiding the complications associated with spontaneous decay seen in the atomic case.